OpenStudy (anonymous):

Find x and y using addition method 5x – 2y = 26 2x – 9y = 35

OpenStudy (anonymous):

Its the elimination method. Ok, lets get rid of the x first. -2(5x-2y=26) <-- I multiplied them so the x would be 0 when subtracted 5(2x-9t=35) -10x+4y=-52 10x-45y=175 -41y=123 y=-3 plug the y=-3 in either equation to solve for x.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I got 4 for x